Status Report

ISS On-Orbit Status 6 Sep 2002

By SpaceRef Editor
September 6, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below. A really busy day today!

Kudos went up to the station residents this morning for an excellent job on yesterday’s SSRMS/Robotics operations. [Checkout and calibration of the POA (payload/ORU accommodation) and the arm’s LEE (latching end effector) went very well, as did the fast grapple and release of the PDGF2 (power and data grapple fixture #2). Two failure messages received during the grapple were false and are understood. An excellent video survey of the MT (mobile transporter) and MBS (mobile base system) was received. The checkout of the PDGF2, the last PDGF to be tested, was conducted subsequently from the ground and was nominal. Korzun and Whitson got good "hands-on" training for 9A and S1 truss transfer operations. The next Robotics Day ops, on 9/12 (Thursday), will feature a LEE camera survey of the nadir ACBM (active common berthing system) on the Node.]

CDR Valery Korzun and FE-2 Sergei Treschev completed the Russian crew health assessment "Profilaktika", part 2, today doing the blood tests with the Reflotron 5 blood analyzer to measure the lactate and creatine kinase levels in their post-exercise blood samples. [Instrument readings were recorded for calldown to MCC-M. Later, Korzun copied accumulated Profilaktika gas analyzer and ECG data to the MEC (medical equipment computer) for downlink via OCA.]

Afterwards, Korzun and Treschev assembled the necessary cables to connect the Service Module’s (SM) Regul interface unit (BSR) to the Regul’s TLFZh channel. The cabling was put in place but not yet hooked up, awaiting connection to two additional units (CA325-I, CA325-II). [Regul OS is the Russian radio control and communications system (RSUS), equivalent to the U.S. S-band system. It is the nominal uplink path for all Russian commanding, and it handles two-way voice comm, digital/command/program data, as well as telemetry via RGS (Russian ground sites).]

FE-2 Treschev performed another atmospheric microbial air sampling using the Russian MO-21 "Ecosphera" air sampler and incubation equipment. [It determines microbial contamination of the ISS atmosphere, specifically the total bacterial and fungal microflora counts and microflora composition according to morphologic criteria of microorganism colonies.]

FE-1 Peggy Whitson continued her science work on the MSG (microgravity science glovebox). Video downlink of yesterday’s SUBSA-08 run indicates good results: the semiconductor alloy melted about 2mm into the dopant seed, as required by the PI (principal investigator). Today Peggy cleaned the MSG thermal chamber in preparation for tomorrow’s sample run (SUBSA09). After SUBSA09, there remain two more runs, after which the MSG will be extensively cleaned and then reconfigured for the Pore Formation and Mobility Investigation (PMFI) during the 9A stage. [Yesterday, when removing sample SUBSA-08 from the thermal chamber, Peggy encountered some debris in the MSG. It is believed to be from SUBSA-07, first plated onto the chamber during the 07 run, then loosened by Wednesday’s 08 thermal cycling. SUBSA work can continue with basic cleaning of the work volume and/or furnace as required.]

Whitson terminated discharge of the second EVA battery (#2024). As result of the recent troubleshooting the ground has determined the discharge function of the BSA (battery stowage assembly) to be working now without fault.

Sergei Treschev continued the barcode labeling of Service Module (SM) panels started by the Expedition 4 crew and continued by Expedition 5 on 7/15. Original number of panels requiring labeling was 186. [The adhesive-backed barcode labels are applied to the backside of the panels, after a suitable spot has been cleaned, in a location convenient for reading with the BCR (bar code reader).]

Valery Korzun completed the routine inspections of the BRPK-1 water condensate separator and the SKV-2 air conditioner fan. He also did the daily SOSh life support system maintenance, while Peggy Whitson performed the daily Lab payload status checkup, and Sergei Treschev the IMS inventory up-date file preparation.

Today’s major scheduled crew activity was the removal and replacement of a critical RPCM (remote power controller module), LA1B-E, in the Lab (starboard overhead) with one of two onboard spares. This required a lengthy and complex sequence of steps, many by ground command. [Necessary power-downs included UOP6 (utility outlet panel #6), two SSC computers and Lab window heaters. The two GNC MDMs (guidance, navigation & control computers) had to be swapped, after attitude control authority was handed over to the Russian segment (RS). After the swap, the new GNC (#2)had to be recovered in Primary mode, including power-cycling the CMGs, (control moment gyros), RGAs (rate gyro assemblies) and GPS-2 (global position satellite system #1). GPS-1 was powered down, as was the RPCM’s DDCU (dc-to-dc converter unit). Attitude control was then returned to the US segment. The actual replacement of the RPCM unit required physical removal of a shear panel and rotation of Peggy’s TeSS (temporary sleep station) rack. After the replacement, TeSS rotation and other multiple procedures had to be reversed, with power-ups again under ground command.]

MCC-H also uplinked a PPL (pre-positioned load) to the GNC MDM for a CMG data dump, and reconfigured the GPS system for the upcoming new LVLH (local vertical/local horizontal) attitude.

The change from sun-oriented XPOP to LVLH (station "belly" always facing down) is scheduled for today at 4:00 pm EDT, when the solar Beta angle has decreased below -10 degrees.

The planned reconfiguration of the onboard Ethernet crew support network BVS in the Russian segment was postponed because of a missing cable. The required cable will be brought up by Progress 9P, after which the final reconfiguration of the onboard network, connecting Russian and U.S. computers, will be finished. 9P will also deliver a new Russian "Wiener" laptop, and the best location for the OpsLAN computers in the SM is currently being discussed by specialists from both sides.

Humidity levels in the station have been on the increase since TsUP switched from SM air conditioner SKV-1 to SKV-2 at end-August. The inequality of the two units, built by different manufacturers, is thought to be due to their different performance characteristics. These are currently being studied by specialists. [For now, the strategy is to let SKV-2 operate most of the time until humidity (H2O partial pressure) has risen to 11 mmHg, then run SKV-1 until ppH2O has decreased to 9.5 mmHg, and then return to SKV-2 alone. After the analysis, MCC-M may recommend another strategy.]

CDRA (carbon dioxide removal assembly) troubleshooting using vacuum evacuation determined that the check valve of adsorb/desorb bed #1 apparently is not leaking as originally thought. To verify this conclusion, the test will be repeated using the air saving pump which provides a greater volume of air to be pumped through the bed than using vacuum. If the valve is indeed found to be OK, the leak may be in a fitting or connection. More discussion to follow.

A procedure is being developed to work around the broken grounding strap of ER2 ARIS (EXPRESS rack 2/active rack isolation system). The work-around involves adapting existing standard straps. Once installed by the crew, it will allow recalibration of the ARIS after its rotation, but it will not allow science operations. Ground spares will be brought up on 9A, and the procedure for their installation will then be available.

The ISS reboost by Progress 8P is still set for 9/11 (Wednesday) at 8:31pm EDT. The burn will use up all remaining Progress propellants and is estimated to yield about 1.8 m/sec (6 ft/sec) delta-V.

The launch of Progress 9P (M1-9) has now slipped from 9/20 to 9/25 (11:58 am EDT). The vehicle has a problem with its Kurs rendezvous radio system and requires a new antenna, which left Moscow today for Baikonur. [The Progress’ Kurs system will be tested on 9/27 at a distance from ISS of about 30 km, and docking at the SM aft end will follow on 9/29 (approx. 2:30 pm EDT). Progress 8P will separate from the aft end on 9/24. The delay will not affect the planned launch of 9A/STS-112 on 10/2, unless 9P should slip beyond 9/28, since it is preferred to have Progress firmly docked before begin of Shuttle tanking on the pad.]

Yesterday’s target for the Russian Diatomeya ocean surveillance program, looking for bioproductive regions, was the area around Oahu, Hawaii.

An interactive TV media downlink for the entire crew is being scheduled for a PAO event with attendees of "German Space Day" at the DLR German Aerospace Center near Cologne, Germany, on 9/8 (Sunday). Expected participants are Edelgard Bulmahn (German Minister for Research and Education), ESA astronauts Gerhard Thiele & Christer Fuglesang, and three students from Bremen.

Today’s targets for the U.S. CEO (crew earth observations program) were Lahore, Pakistan (monsoon cloudiness is slowly retreating eastwards. Nadir pass over this major city. ESC [electronic still camera]. Crew to try for city margins), Alexandria, Egypt (nadir pass. ESC), Cairo, Egypt (nadir pass. ESC. Pyramids lie on the west side of the base of the Nile delta), Nile River Delta (nadir pass. ESC. Detail of expansion of agricultural lands is of great interest, especially along desert margins on either side of the Suez Canal), Western Mediterranean Dust (small low pressure system over southern France was kicking up dust on the coast of Algeria), Detroit, Michigan (nadir pass. ESC. Crew to try for city margins), Angolan Biomass Burning (crew to look left of track), Johannesburg, South Africa (oblique views contextual shots requested of this center point of a line of cities), and Central Amazon basin (unusually cloud free. Detailed mapping pass of rivers and new roads requested. Left and right [for sunglint opportunity]) of track).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:44 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (16-amp mode), on backup pump. Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed). U.S. CDRA CO2 scrubber on command override. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment: Pressure (mmHg) — 757, temperature (deg C) — 27.2, ppO2 (mmHg) — 145.5, ppCO2 (mmHg) — 0.1.
SM Transfer Compartment: Pressure (mmHg) — 752, temperature (deg C) — 21.2.
FGB Cabin: Pressure (mmHg) — 752, temperature (deg C) — 21.7.
Node: Pressure (mmHg) — 748.98, temperature (deg C) — 24.9 (shell); ppO2 (mmHg) — 161.9*; ppCO2 (mmHg) — 4.7*.
U.S. Lab: Pressure (mmHg) — 751.38, temperature (deg C) — 24.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 750.98, temperature (deg C) — 27.8; shell heater temp (deg C) — 25.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1: Shell heater temp (deg C) — 24.3
PMA-2: Shell heater temp (deg C) — 15.5

(*Note: Partial pressures ppO2 and ppCO2 in Node are obtained from the recently restored MCA [major constituent analyzer] and are not considered reliable until purge of water vapor within sample lines has been completed.

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and BGA 4B in AutoTrack mode (solar-tracking).
SM batteries: Battery #1 is offline, battery #7 is in "Capacity Restoration" mode (ROM); all other batteries (6) are in "Partial Charge" mode. Data are static.
FGB batteries: Battery #4 is in "Capacity Restoration" mode (ROM), battery #6 is cycling; all other batteries (4) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is Off; SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-2 MDM is prime; GNC-1 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is Off (cold backup); PL-2 MDM is operating as primary.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; #3 dropped out on 8/29.

Attitude Source:
3 CMGs on-line.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF1 (mobile base system/power & data grapple fixture 1) and PDGF2, with Keep Alive power on both strings. SSRMS based on PFDGF1 (LEE A, derigidized)
MBS: Keep Alive power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 8:06am EDT [= epoch]):
Mean altitude — 389.4 km
Apogee — 399.7 km
Perigee — 379.1 km
Period — 92.3 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0015205
Orbits per 24-hr. day — 15.59
Solar Beta Angle — -11.4 deg (magnitude decreasing)
Altitude decrease — 200 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ’98) — 21671
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: -0.5 deg, pitch: -5.7 deg., roll: 0 deg]). Will change to LVLH today at 4:00 pm EDT, when solar Beta drops below -10 deg.

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.